OpenStudy (anonymous):

The ordered pair (4,2) is a solution to the system of inequalities below. y>-3x+2 y<2x+1 True False

OpenStudy (anonymous):

put wha x and y is in both cases

OpenStudy (anonymous):

2>-3*4+2=-12+2=-10 meaning 2>-10 true

OpenStudy (anonymous):

nw test 2nd one